determine Laplace transform of a-d
(a) f(1) = (1 - 4)u(t - 2) (b) g(t) = 2e-4eu(t - 1) (c) h(t) = 5 cos(2t - 1)u(t) (d) p(t) = 6[u(t - 2) - ut - 4)]
Answer #1


FIS) 5 - 2025 v Laplace trasform Srls = 10 fev erot at? formula a f 4+1 = (t-4) U(+-2) but the value in equn FIS) = SW (t-4)-at ult-1) 6 geH= = 2e using formula of Laplace dt 915) - Lo quest - J^2 eftult-1) estali Ietlo U16-1)- other Ult-indt = 21°using table of Laplace transform ILI Cas wot S s²two? Sin not a LT Wo 5²+ug? htl=5 Cosat ult) + 0.087 Sinat ult) taking taplaPltl= 2 <ta4 6 1 otherwise 4 dt PIS) = 16 6e-st 2 4 = est 6 2 نی -s -25 56 e te [ PIS) = 6 -s P(S) - 6 & [e-es Ang